solve the system of equations: y=3x-10 and x=12-4y
Answer by checkley77(12844)

y=3x-10
x=12-4y
x=12-4(3x-10)
x=12-12x+40
x+12x=52
13x=52
x=52/13
x=4 ans.
y=3*4-10
y=12-10
y=2 ans.
Proof:
4=12-4*2
4=12-8
4=4